The pressure on 6dm3 volume of gas is 200kPa. What will be the volumn if the pressure is doubled, keeping the temperature constant


1
Expert's answer
2021-08-16T08:36:02-0400

"V_1= 6 \;dm^3 \\

p_1= 200 \;kPa \\

p_2 = 2 \times 200 = 400 \;kPa \\

p_1V_1=p_2V_2 \\

V_2 = \frac{p_1V_1}{p_2} \\

= \frac{200 \times 6}{400} = 3 \;dm^3"

Answer: 3 dm3